Circular No. 7100

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 990123 C. W. Akerlof and T. A. McKay, University of Michigan, report on

behalf of the ROTSE collaboration (Michigan/Los Alamos National Laboratory/Lawrence Livermore National Laboratory): “We observed the error box of GRB 990123 provided by the BACODINE Burst Position Notice (dated 23-Jan-99 09:46:59) using the ROTSE-I telephoto camera array located at Los Alamos. The first exposure began at 9^{h}47^{m}18^{s}.30 UT (Jan. 23.407851), or 22.18 s after the nominal burst trigger time. A rapidly fading object was discovered at α = 15^{h}25^{m}30^{s}.1, δ = +44^{o}46′00′′ (equinox 2000.0), which is within one-third of a pixel of the optical counterpart reported by Odewahn et al. (IAUC 7094). The lightcurve for this object is relatively complex: the luminosity increases by 3 magnitudes between the first and second exposures. Exposure times and estimated V magnitudes for the first six exposures are as follows: Jan. 23.407851, 5 s, 11.82; 23.408142, 5, 8.95; 23.408435, 5, 10.08; 23.410851, 75, 13.22; 23.412764, 75, 14.00; 23.414677, 75, 14.53. Note that the ROTSE-I detector system uses an unfiltered broadband CCD, so that magnitude estimates are based on comparisons to catalogue values for nearby stars. Sky-patrol images of the same coordinates taken 133 min earlier showed no evidence of the transient to a limit of at least 2 mag deeper. A more extensive analysis of these data will be available in the near future. The discovery images are posted on the ROTSE Web

page at: http://www.umich.edu/~rotse/gifs/grb990123/990123.gif.”

J. Zhu, J. S. Chen, and H. T. Zhang, Beijing Astronomical Observatory (BAO), report that an observation on Jan. 24.730–24.861 UT with the BAO 0.6-m Schmidt telescope under nonphotometric conditions (two 60min images taken with a BATC i-band filter; central wavelength 666.0 nm, bandwidth 48.0 nm) gives a magnitude of 21.0 ± 0.3 (Jan. 24.818) for the transient + presumed host galaxy, using stars 1 and 2 from GCN 204

(http://gcn.gsfc.nasa.gov/gcn/gcn3/204.gcn3). No visual separation

between the optical transient and host galaxy could be recognized because of the low S/N ratio and bad spatial resolution (1 pixel = 1′′.7; seeing 5′′.6).

The CCD image is available at http://vega.bac.pku.edu.cn/~zj/grb/

grb990123.html. A 40-min exposure on Jan. 25.901 yields a magnitude of 21.3 ± 0.3 (near the detection limit), suggesting that by this time they were only detecting the the coincident galaxy that was found on the digital sky-survey image by Odewahn et al. (IAUC 7094).

1999 January 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green